# Quantum metrics from length functions on quantum groups

### v1

@@ 6. Examples
The aim of the present section is to prove Corollary C, and thereby provide concrete examples of length functions on quantum groups whose associated seminorms give compact quantum metric structures. Our primary focus will therefore be on compact quantum matrix groups [Tim08, Section 6.1], and below we will restrict further to the situation where the fundamental corepresentation is equivalent to its conjugate. In this situation, every irreducible corepresentation $\alpha$ is equivalent to a subrepresentation of some iterated tensor power of the fundamental unitary corepresentation, and setting $\ell(\alpha)$ equal to the first tensor power containing $\alpha$ defines a length function on the quantum group in question. We first prove the following result regarding length functions of this type.
@@
Let ${\mathbb{G}}$ be a compact matrix quantum group with fundamental unitary corepresentation $v$. Suppose, moreover, that $v$ is equivalent to its conjugate and that the associated length function $\ell\colon\operatorname{Irred}({\mathbb{G}})\to\mathbb{N}_{0}$ is injective. If the set of structure constants $\{N_{\alpha,\beta}^{\gamma}\mid\alpha,\beta,\gamma\in\operatorname{Irred}({\mathbb{G}})\}$ is bounded, then $(F({\mathbb{G}}),\ell^{2}(\operatorname{Irred}({\mathbb{G}}),\tilde{D}_{\ell})$ is a spectral metric space.
@@
Denote the image of $\ell$ as $\{n_{0},n_{1},n_{2},\ldots\}\subseteq{\mathbb{N}}_{0}$ (listed in increasing order) and label the elements in $\operatorname{Irred}({\mathbb{G}})$ accordingly as $\{\alpha_{0},\alpha_{1},\alpha_{2},\ldots\}$ so that $\ell(\alpha_{k})=n_{k}$ for all $k\in{\mathbb{N}}_{0}$. For $n\in{\mathbb{N}}_{0}$, denote by $A_{n}$ the finite-dimensional subspace $\mathrm{span}_{\mathbb{C}}\{\alpha\in\operatorname{Irred}({\mathbb{G}})\mid\ell(\alpha)\leqslant n\}\subseteq F({\mathbb{G}})$. In Section 3.1, we saw that the fusion algebra together with its natural trace and the filtration $(A_{n})_{n\in\mathbb{N}_{0}}$ falls within the class studied by Ozawa and Rieffel in [OR05], and we now wish to use their Main Theorem 1.2 to show our result. Denote by $P_{n}\in{\mathbb{B}}(\ell^{2}(\operatorname{Irred}({\mathbb{G}})))$ the projection onto the subspace spanned by irreducible corepresentations of length exactly $n$, with the convention that $P_{n}=0$ if there are no such corepresentations. Decomposing an element $a\in F({\mathbb{G}})$, accordingly, as a finite sum $\sum_{k=0}^{\infty}a_{k}$ as in [OR05, Main Theorem 1.2], we have $a_{k}=0$ if $k\notin\{n_{0},n_{1},n_{2},\dots\}$ and for $k\in\{n_{0},n_{1},n_{2},\dots\}$, $a_{k}$ takes the form $z_{k}\alpha_{k}$ for some $z_{k}\in{\mathbb{C}}.$ To verify the conditions of [OR05, Main Theorem 1.2], we must provide a constant $C>0$ such that for all $a\in F({\mathbb{G}})$ and all $k,m,n\in{\mathbb{N}}_{0}$ it holds that
@@
from which it follows that $\ell(u^{n})=n$ and that that the structure constants are uniformly bounded by 1. The assumptions in Lemma 6.1 are therefore satisfied and we conclude that $(F(SU(2)),\ell^{2}(\operatorname{Irred}(SU(2))),\tilde{D}_{\ell})$ is a spectral metric space. Since $SU(2)$ is coamenable and of Kac type, an application of Corollary 3.7 and Theorem A shows that $(C(SU(2)),L_{\ell})$ is a compact quantum metric space.
Turning to $O_{2}^{+}$, by [Ban96] its irreducible corepresentations may also be labeled as $\{u^{n}\}_{n\in{\mathbb{N}}_{0}}$ with $u^{0}$ being the trivial corepresentation, $u^{1}$ being the self-conjugate fundamental unitary corepresentation and with fusion rules described by (6.2). The associated length function therefore also satisfies $\ell(u^{n})=n$. Since $O_{2}^{+}$ is coamenable and of Kac type and since we have already proved that $(C(SU(2)),L_{\ell})$ is a compact quantum metric space, an application of Corollary B yields the corresponding result for $O_{2}^{+}$.
@@
and it therefore follows that $\ell(u^{n})=n$ and that the structure constants are uniformly bounded by $1$. By Lemma 6.1, we therefore conclude that $(F(SO(3)),\ell^{2}(\operatorname{Irred}(SO(3))),\tilde{L}_{\ell})$ is a spectral metric space, and combining Corollary 3.7 and Theorem A we conclude that $(C(SO(3)),L_{\ell})$ is a compact quantum metric space. To finish the proof, we note that it was shown in [Ban99b] that the corepresentation theory of $S_{4}^{+}$ and $SO(3)$ are identical (in the same way as was the case for $SU(2)$ and $O_{2}^{+}$) and since $S_{4}^{+}$ is both coamenable and of Kac type, Corollary B implies that also $(C(S_{4}^{+}),L_{\ell})$ is a compact quantum metric space. ∎
@@
We end with an aside regarding finite-dimensional truncations, a subject which has recently attracted a lot of attention [GEvS23, van21, Cv21, Cv22, LvS24, Lei24]). Rieffel’s recent paper [Rie23] (see also [Lei24]) provides a systematic framework for coactions of quantum groups, within which such finite-dimensional approximations can be obtained. The examples studied above are compatible with the main result (Theorem 6.1) in [Rie23]; indeed, our context is also restricted to the realm of coamenable compact quantum groups, and in Lemma 4.5 we have proven that slip-norms arising from length functions are left invariant. Note, in this respect, that the conventions in [Rie23] are different than those in the present paper, in that the left coactions of [Rie23] are what would be called right coactions in our setup. Paraphrasing [Rie23, Theorem 6.1] in the context of (our) left coactions, means that the necessary invariance property of the Lip-norm in question is exactly the one covered by Lemma 4.5. Hence, for ergodic coactions of the examples covered by Corollary C, [Rie23] provides Lip-norms on the algebra acted upon (this construction is originally due to Li [Li09]) together with a natural sequence of finite-dimensional subspaces converging to the total space in the quantum Gromov-Hausdorff distance [Rie04b].

### v3

@@ 6. Examples
The aim of the present section is to prove Corollary C, and thereby provide concrete examples of length functions on quantum groups whose associated seminorms give compact quantum metric structures. Our primary focus will therefore be on compact matrix quantum groups [Timmermann-book, Section 6.1], and below we will restrict further to the situation where the fundamental corepresentation is equivalent to its conjugate. In this situation, every irreducible corepresentation $\alpha$ is equivalent to a subrepresentation of some iterated tensor power of the fundamental unitary corepresentation, and setting $\ell(\alpha)$ equal to the first tensor power containing $\alpha$ defines a length function on the quantum group in question. We first prove the following result regarding length functions of this type.
@@
[…]
@@
Denote the image of $\ell$ as $\{n_{0},n_{1},n_{2},\ldots\}\subseteq{\mathbb{N}}_{0}$ (listed in increasing order) and label the elements in $\operatorname{Irred}({\mathbb{G}})$ accordingly as $\{\alpha_{0},\alpha_{1},\alpha_{2},\ldots\}$ so that $\ell(\alpha_{k})=n_{k}$ for all $k\in{\mathbb{N}}_{0}$. For $n\in{\mathbb{N}}_{0}$, denote by $A_{n}$ the finite-dimensional subspace $\mathrm{span}_{\mathbb{C}}\{\alpha\in\operatorname{Irred}({\mathbb{G}})\mid\ell(\alpha)\leqslant n\}\subseteq F({\mathbb{G}})$. In Section 3.1, we saw that the fusion algebra together with its natural trace and the filtration $(A_{n})_{n\in\mathbb{N}_{0}}$ falls within the class studied by Ozawa and Rieffel in [OzawaRieffel2005], and we now wish to use their Main Theorem 1.2 to show our result. Denote by $P_{n}\in{\mathbb{B}}(\ell^{2}(\operatorname{Irred}({\mathbb{G}})))$ the projection onto the subspace spanned by irreducible corepresentations of length exactly $n$, with the convention that $P_{n}=0$ if there are no such corepresentations. Decomposing an element $a\in F({\mathbb{G}})$, accordingly, as a finite sum $\sum_{k=0}^{\infty}a_{k}$ as in [OzawaRieffel2005, Main Theorem 1.2], we have $a_{k}=0$ if $k\notin\{n_{0},n_{1},n_{2},\dots\}$ and for $k\in\{n_{0},n_{1},n_{2},\dots\}$, $a_{k}$ takes the form $z_{k}\alpha_{k}$ for some $z_{k}\in{\mathbb{C}}.$ To verify the conditions of [OzawaRieffel2005, Main Theorem 1.2], we must provide a constant $C>0$ such that for all $a\in F({\mathbb{G}})$ and all $k,m,n\in{\mathbb{N}}_{0}$ it holds that
@@
[…]
Turning to $O_{2}^{+}$, by [Banica-On+-rep] its irreducible corepresentations may also be labeled as $\{u^{n}\}_{n\in{\mathbb{N}}_{0}}$ with $u^{0}$ being the trivial corepresentation, $u^{1}$ being the self-conjugate fundamental unitary corepresentation and with fusion rules described by (6.2). The associated length function therefore also satisfies $\ell(u^{n})=n$. Since $O_{2}^{+}$ is coamenable and of Kac type and since we have already proved that $(C(SU(2)),L_{\ell})$ is a compact quantum metric space, an application of Corollary B yields the corresponding result for $O_{2}^{+}$.
@@
[…]
@@
We end with an aside regarding finite-dimensional truncations, a subject which has recently attracted a lot of attention [walter-estrada-cp-grps, walter:GH-convergence, walter-connes:truncations, CvS:Truncations, walter-tori, Leimbach:PWT]). Rieffel’s recent paper [Rie:truncations] (see also [Leimbach:PWT]) provides a systematic framework for coactions of quantum groups, within which such finite-dimensional approximations can be obtained. The examples studied above are compatible with the main result (Theorem 6.1) in [Rie:truncations]; indeed, our context is also restricted to the realm of coamenable compact quantum groups, and in Lemma 4.5 we have proven that slip-norms arising from length functions are left invariant. Note, in this respect, that the conventions in [Rie:truncations] are different than those in the present paper, in that the left coactions of [Rie:truncations] are what would be called right coactions in our setup. Paraphrasing [Rie:truncations, Theorem 6.1] in the context of (our) left coactions, means that the necessary invariance property of the Lip-norm in question is exactly the one covered by Lemma 4.5. Hence, for ergodic coactions of the examples covered by Corollary C, [Rie:truncations] provides Lip-norms on the algebra acted upon (this construction is originally due to Li [Li:ECQ]) together with a natural sequence of finite-dimensional subspaces converging to the total space in the quantum Gromov-Hausdorff distance [Rie:GHD].